OpenStudy (anonymous):

write an equation of the line containing the given point & perpendicular to the given line. (0,8); 5x+8y=7

OpenStudy (anonymous):

use the point-slope formula again this time acknowledging the fact that the slopes of perpendicular lines are negative reciprocals of each other (i.e. m1=2, m2=-1/2). You will obtain y=8-(8/5)x